SHOW LIST OF ACCOUNTS LINK TO TEXT Compute the total net liability to be reported on the December 31 balance sheet for (1) The interest-bearing note (2) The zero-interest-bearing note Click if you would like to Show Work for this question: Open Show Work SHOW LIST OF ACCOUNTS LINK TO TEXT O Type here to search

Exercise 13-2 The following are selected 2017 transactions of Larkspur Corporation Sept. 1 Purchased inventory from Encino Company on account for $58,600. Larkspur records purchases gross and uses a periodic inventory system Oct. 1 Oct. 1 Issued a $58,600, 12-month, 8% note to Encino in payment of account. Borrowed $58,600 from the Shore Bank by signing a 12-month, zero-interest-bearing $63,720 note. reauired, select No Entry for the account titles and enter 0 for the am
Answer #1

Answers

  • All working forms part of the answer
  • Requirement 1

A

Amount of Interest bearing note

$58,600

B

Interest rate

8%

C= A x B

Annual 12 month interest

$4,688

Period till 31 Dec

1 Oct to 31 Dec

= 3 months

D = C x 3/12

Interest Payable accrued for 3 months

$1,172

E = A+D

Net Liability to be reported

$59,772 = Answer

  • Requirement 2

A

Zero bearing note amount

$63,720

B

Amount borrowed

$58,600

C = A - B

12 month interest = Discount on issuance

$5,120

Period till 31 Dec

1 Oct to 31 Dec

= 3 months

D = C x 3/12

Interest payable for 3 months

$1,280

E = C - D

Net Discount

$3,840

F = A - E

Net Liability to be reported

$59,880 = Answer